In your chemistry lab, you have a bottle of 2% hydrochloric acid solution and a bottle of 6% hydrochloric solution. You need 100 millimeters of 3% hydrochloric acid solution for an experiment. How many millimeters of each solution do you need to mix together?
Answer by stanbon(75887)

In your chemistry lab, you have a bottle of 2% hydrochloric acid solution and a bottle of 6% hydrochloric solution. You need 100 millimeters of 3% hydrochloric acid solution for an experiment. How many millimeters of each solution do you need to mix together?
----------
Equation:
acid + acid = acid
0.02x + 0.06(100-x) = 0.03*100
------------------
2x + 6*100 - 6x = 3*100
-4x = -3*100
-------
x = (3/4)100
-------
x = 75 ml (amt. of 2% solution needed)
-----
100-x = 25 ml (amt. of 6% solution needed)
=============================================
Cheers,
stan H.
--------------------